OpenStudy (stuck-help):

Given: triangle QRS~= triangle TUV,QS=3v+2and TV-7v-6, find the length of QS and TV

OpenStudy (johnweldon1993):

Hint* If they are congruent, they are equal So the 2 sides they are comparing are congruent \[\large 3v + 2 = 7v - 6\] Just solve for 'v' and then plug that into both equations to find the length of each

OpenStudy (stuck-help):



OpenStudy (stuck-help):

@johnweldon1993

OpenStudy (johnweldon1993):

Right so if V= 2 Length of QS = 3v + 2 = 3(2) + 2 = 6 + 2 = 8 And Length of TV = 7v - 6 = 7(2) - 6 = 14 - 6 = 8 Makes sense because remember since the triangles are congruent, they will have the same side length